# Evaluating the Impacts of Swapping on the US Decennial Censusthanks: This work was supported in part by Sloan Foundation grant G-2017-9890, a Google Research Scholar award, and the Center for Research on Computation and Society (CRCS), Harvard

### v1

@@ Abstract
To meet its dual burdens of providing useful statistics and ensuring privacy of individual respondents, the US Census Bureau has for decades introduced some form of “noise” into published statistics, initially through a method known as “swapping” (1990–2010), and then, for the first time in 2020, via an algorithm ensuring a form of Differential Privacy. While the TopDown algorithm used in 2020 has been made public, no implementation of swapping has been released, in part to preserve the confidentiality of respondent data. The Bureau has not published (even a synthetic) “original” dataset and its swapped version, and it has kept secret many details of the swapping methodology deployed. It is therefore difficult to evaluate the effects of swapping, and to compare swapping to other privacy technologies. To address these difficulties we describe and implement a parameterized swapping algorithm based on Census publications and court documents, and informal interviews with Census employees. With this implementation, we characterize the impacts of swapping on a range of statistical quantities of interest. We provide intuition for the types of shifts induced by swapping and compare against techniques that use differential privacy. We find that even when swapping and differential privacy introduce errors of a similar magnitude, the direction in which statistics are biased need not be the same across the two techniques. More broadly, our implementation provides researchers with the tools to analyze and potentially correct for the impacts of disclosure avoidance systems on the quantities they study.
@@ 1 Introduction
Census statistics are used for a wide variety of planning, administrative, and redistricting purposes. In order to meet its dual burdens of providing useful statistics and protecting the privacy of individual respondents, the US Census Bureau has for decades introduced noise into published statistics. From 1990–2010, the decennial census used a form of “swapping” (Dalenius and Reiss, 1982; McKenna, 2019). Because the swapping method used by the Census provides no provable guarantees, the method itself is kept confidential in an effort to maintain respondent privacy.
@@
At a high level, swapping works by choosing households that are at a particularly high disclosure risk—say, a household that is racially unique in its neighborhood—and “swapping” each such household with another household in a different region. Effectively, swapping works by “hiding” racially unique households in areas where they may be less identifiable. The precise instantiation of this framework used by the Bureau is kept secret, and indeed, we do not know whether any single algorithm was used across all states in the 1990–2010 decennial censuses.
@@
Prior to 2020, although it was generally known that swapping was deployed, no details were available and consumers of Census statistics tended to treat these data as ground truth. This became impossible with the announcement of the deployment of differential privacy and the publication of demonstration data and, eventually, the TopDown algorithm itself. In a nutshell, consumers were forced to confront the “ground truth” fiction. Each DAS, swapping and TopDown, has both privacy properties and impacts on data analysis. We know that the Bureau chose TopDown because it can provide better privacy protection than that obtained with the level of swapping used in the 2010 census. In this work, we turn to the impact of these disclosure avoidance systems on data analysis.
@@
Studying the impacts of disclosure avoidance systems is challenging due to a lack of both data and information. In the case of TopDown, the Census Bureau has provided both the details of the algorithm and a series of demonstration data releases to facilitate research (see Section 2 for details). No analogous releases exist for swapping. In this work, we facilitate research into the effects of swapping through a careful implementation of a swapping algorithm based on information released by the Census Bureau (see Section 3); specifically, we drawn from a variety of publications, presentations, and statements by Bureau employees as well as court documents to create an approximation to the Bureau’s practices. For details that remain unknown, we conduct robustness checks to verify that our conclusions hold across a range of reasonable choices.
@@
We use this implementation to study a variety of the impacts that swapping has on the Census data. In Section 4, to investigate the basic statistical properties of swapping, we measure the errors that swapping introduces to population counts by race and compare them to those of TopDown and a simple variant of TopDown created by Cohen et al. (2021) called ToyDown, at a specific setting and apportioning of the privacy parameter $\varepsilon$. We further examine detailed properties of swapping with respect to household size, mixed-race households, and racial entropy across geographies. Surprisingly, despite the fact that swapping targets households that are racially unique in their current location and “moves” them elsewhere, we find that swapping increases racial entropy. We provide a detailed analysis of why this is the case.
@@
We next consider the impact of swapping on downstream uses of Census data for both public policy and social science (Section 5). Prior work has found that differential privacy biases statistics derived from ecological regression (ER) analyses, which are used in cases involving the Voting Rights Act (Cohen et al., 2021). We find that the same is true for swapping, but crucially, swapping appears to bias these statistics in the opposite direction. We again provide intuition for why this might be the case. Thus, viewing swapping as a similarly noisy predecessor to differential privacy may obscure key differences in the direction of effects of that noise.
@@
We further compare swapping to TopDown in terms of their impact on social science applications, where recent work has shown that the noise introduced by the TopDown algorithm can lead to substantial variance in downstream analyses (Mueller and Santos-Lozada, 2022; Hauer and Santos-Lozada, 2021; Ruggles et al., 2019; Santos-Lozada et al., 2020; Winkler et al., 2021). In particular, Mueller and Santos-Lozada (2022) show that TopDown can introduce large errors in rural areas. We find that the same is true for swapping. We conclude by proposing a more general framework for using our methods to estimate the impacts of swapping on downstream statistics of interest.
@@ 2 Related Work
Our focus in this work is on the Decennial Census. Disclosure avoidance is also used in the American Community Survey (ACS), which collects more detailed information on a sample of the population (see Zayatz et al. (2009)). The Census Bureau uses a variety of disclosure avoidance techniques to release ACS data, including swapping (Poehler, 2022, Section 13.8), Census Bureau (2022), Rodríguez (2023).
@@ 3 Swapping Implementation
The swap rate is not public, but we have some information about reasonable ranges. Bailie et al. (2023) claim that the swap rate is 2–4%. John Abowd, the former Chief Scientist at the U.S. Census Bureau, states that the swap rates that the Bureau considered when studying the feasibility of using swapping in 2020 were higher than the swap rates used in 2010 (Abowd, 2021). An example of the swap rates that the Bureau considered for 2020 appears in Hawes and Rodríguez (2021) (also written by Bureau officials), in which the authors present swap rates in the range 5-50%. Based on this information, we use 2% as a reasonable low swap rate and and 10% as a reasonable high swap rate. In some places, we also present results for higher swap rates.
@@
Documentation on matching variables is inconsistent. According to Zayatz (2002), “The key variables were number of people in the household of each race by Hispanic/NonHispanic by age group (<18,18+), number of units in building, rent/value of home, and tenure.” Griffin et al. (1989) states that the planned swapping methodology matches on “number of persons in household, population characteristics of race, Hispanic origin and age, and on housing characteristics of units at building, rent/value and tenure.” If true, however, this would guarantee that block-level counts by race and ethnicity are invariant to swapping, which goes against assumptions commonly held in the literature (e.g., (Christ et al., 2022; Kenny et al., 2023b)). Moreover, Steel and Zayatz (2003) appear to contradict this in their evaluation of the impacts of swapping, writing “Our procedure targeted households with unique characteristics, and this increased the number of zero cells (for these characteristics) because it tended to draw people from cells with values 1, 2, and 3 (presumably where the entire contribution to the cell is from one household and hence selected for high disclosure risk) and swap them into cells with larger values.”
@@ 4.1 Setup and notation
By running ToyDown, the simplified version of TopDown implemented by (Cohen et al., 2021), on our synthetic data, we can also produce $\mathbf{X}_{\text{ToyDown}}$. To produce this database, we run ToyDown with $\epsilon=3.26$ and an equal split of privacy budget across geographic levels. This value of $\epsilon$ was chosen by calibrating the variance of ToyDown against the variance of TopDown in demonstration data (see Section 4.3 for details). Technically, our implementation of ToyDown differs from the one used in (Cohen et al., 2021) because we do not treat Hispanic as a race category to be consistent with our swapping algorithm.
@@
Figure 1: Various versions of the Census data and the relationships between them. Note that there are two releases of $\mathbf{X}_{\text{TopDown}}$, which we call $\mathbf{X}_{\text{TopDown2021}}$ and $\mathbf{X}_{\text{TopDown2023}}$. We also produce multiple versions of $\mathbf{X}_{\text{Swapped}}$: $\mathbf{X}_{\text{Swapped2}}$ is produced by swapping at a 2% swap rate and $\mathbf{X}_{\text{Swapped10}}$ is produced by swapping at a 10% swap rate. The red arrows show the comparisons made in the panels in Figures 2 and 3. The blue arrow shows the comparison of interest (which we cannot directly observe).
@@ 4.2 Quantifying the Errors Introduced by Swapping
We begin by studying the bias of swapping on race counts. It is important to remember that the state-wide average bias of the count of any race is zero, since swapping (unlike TopDown) does not change the statewide counts of each race. Therefore, we study the error in the individual counts at lower geographic levels. These errors will average to zero, but we are interested in their magnitudes.
@@
Figure 2 shows $\textsf{Error}_{\gamma}^{r}(\mathbf{X}_{\text{Synthetic}},\mathbf{X}_{\text{Swapped2}})$, $\textsf{Error}_{\gamma}^{r}(\mathbf{X}_{\text{Synthetic}},\mathbf{X}_{\text{Swapped10}})$, $\textsf{Error}_{\gamma}^{r}(\mathbf{X}_{\text{Synthetic}},\mathbf{X}_{\text{ToyDown}})$, and $\textsf{Error}_{\gamma}^{r}(\mathbf{X}_{\text{Released}},\mathbf{X}_{\text{TopDown2021}})$ at the county level in Alabama. As noted earlier, the errors for swapping (at both 2% and 10%) are both mean zero. [^4] The swap errors do not necessarily have median zero; the blue lines in the boxplots are the medians. We observe that errors due to swapping (Panels A and B) are generally greater than errors due to ToyDown (Panel C). Also, swapping results in higher errors on large racial groups, while ToyDown distributes errors evenly across racial groups. The picture for TopDown (Panel D) may superimpose the errors due to swapping and TopDown, which is consistent with the fact that both swapping and TopDown stand between $\mathbf{X}_{\text{TopDown}}$ and $\mathbf{X}_{\text{Released}}$.
@@ C Refer to caption Refer to caption D
Figure 2: Plots of the four quantities $\textsf{Error}_{\gamma}^{r}(\mathbf{X}_{\text{Synthetic}},\mathbf{X}_{\text{Swapped2}})$, $\textsf{Error}_{\gamma}^{r}(\mathbf{X}_{\text{Synthetic}},\mathbf{X}_{\text{Swapped10}})$, $\textsf{Error}_{\gamma}^{r}(\mathbf{X}_{\text{Synthetic}},\mathbf{X}_{\text{ToyDown}})$, $\textsf{Error}_{\gamma}^{r}(\mathbf{X}_{\text{Released}},\mathbf{X}_{\text{TopDown2021}})$ for all $\gamma\in\mathcal{C}_{\text{Alabama}}$, where $\mathcal{C}_{\text{Alabama}}$ is the set of all counties in Alabama. To generate the “Average” plot, for each county, the errors for each race were averaged. Note that the scale of the $y$-axes are different in different plots.
@@
as a measure of the effect of swapping on the count of race $r$ in region $\gamma$, normalized by the size of the count. If the count increases by a relatively large amount from $\mathbf{X}_{1}$ to $\mathbf{X}_{2}$, $\textsf{RelativeError}_{\gamma}^{r}(\mathbf{X}_{1},\mathbf{X}_{2})$ will be close to 2. If the count decreases by a relatively large amount, $\textsf{RelativeError}_{\gamma}^{r}(\mathbf{X}_{1},\mathbf{X}_{2})$ will be close to 0. If the count is unchanged, $\textsf{RelativeError}_{\gamma}^{r}(\mathbf{X}_{1},\mathbf{X}_{2})=1$. If both counts are 0, we treat $\textsf{RelativeError}_{\gamma}^{r}(\mathbf{X}_{1},\mathbf{X}_{2})$ as 1. In practice, the way we treat zeros is not very important because the only county-level race counts that are zero are the number of Hawaiian or Pacific Islander people in five counties (before swapping).
Again, our ideal comparison is $\textsf{RelativeError}_{\gamma}^{r}(\mathbf{X}_{\text{CEF}},\mathbf{X}_{\text{Released}})$, but since we do not observe $\mathbf{X}_{\text{CEF}}$, we instead use $\textsf{RelativeError}_{\gamma}^{r}(\mathbf{X}_{\text{Synthetic}},\mathbf{X}_{\text{Swapped}})$ (defined for swapping at a 2% swap rate and a 10% swap rate), as well as $\textsf{RelativeError}_{\gamma}^{r}(\mathbf{X}_{\text{Synthetic}},\mathbf{X}_{\text{ToyDown}})$ and $\textsf{RelativeError}_{\gamma}^{r}(\mathbf{X}_{\text{Released}},\mathbf{X}_{\text{TopDown2021}})$. In Figure 3, we show the relative errors $\textsf{RelativeError}_{\gamma}^{r}(\mathbf{X}_{\text{Synthetic}},\mathbf{X}_{\text{Swapped2}})$, $\textsf{RelativeError}_{\gamma}^{r}(\mathbf{X}_{\text{Synthetic}},\mathbf{X}_{\text{Swapped10}})$, $\textsf{RelativeError}_{\gamma}^{r}(\mathbf{X}_{\text{Synthetic}},\mathbf{X}_{\text{ToyDown}})$, $\textsf{RelativeError}_{\gamma}^{r}(\mathbf{X}_{\text{Released}},\mathbf{X}_{\text{TopDown2021}})$. We observe that 2% swapping has smaller relative errors than ToyDown, but 10% swapping has comparable relative errors to ToyDown. Both swapping and ToyDown exhibit very low relative error on large racial groups and much larger relative error on small racial groups (especially with a 10% swap rate, but often with a 2% swap rate as well). In the case of ToyDown, this is consistent with the fact that ToyDown distributes absolute errors relatively evenly across racial groups, so groups with smaller sizes have larger relative errors. In the case of swapping, it appears that, while swapping has larger absolute errors on larger racial groups, this is more than cancelled out by the effect of the group sizes. The overall effect is that swapping, at least at a 10% swap rate, introduces larger relative errors to smaller groups.
@@
Figure 3: Plots of the four quantities $\textsf{RelativeError}_{\gamma}^{r}(\mathbf{X}_{\text{Synthetic}},\mathbf{X}_{\text{Swapped2}})$, $\textsf{RelativeError}_{\gamma}^{r}(\mathbf{X}_{\text{Synthetic}},\mathbf{X}_{\text{Swapped10}})$, $\textsf{RelativeError}_{\gamma}^{r}(\mathbf{X}_{\text{Synthetic}},\mathbf{X}_{\text{ToyDown}})$, and $\textsf{RelativeError}_{\gamma}^{r}(\mathbf{X}_{\text{Released}},\mathbf{X}_{\text{TopDown2021}})$ for all $\gamma\in\Gamma$, where $\Gamma$ is the set of all counties in Alabama. To generate the “Average” plot, for each county, the relative errors for each race were averaged.

### v2

@@ Abstract
To meet its dual burdens of providing useful statistics and ensuring privacy of individual respondents, the US Census Bureau has for decades introduced some form of “noise” into published statistics. Initially, they used a method known as “swapping” (1990–2010). In 2020, they switched to an algorithm called TopDown that ensures a form of Differential Privacy. While the TopDown algorithm has been made public, no implementation of swapping has been released and many details of the deployed swapping methodology deployed have been kept secret. Further, the Bureau has not published (even a synthetic) “original” dataset and its swapped version. It is therefore difficult to evaluate the effects of swapping, and to compare these effects to those of other privacy technologies. To address these difficulties we describe and implement a parameterized swapping algorithm based on Census publications, court documents, and informal interviews with Census employees. With this implementation, we characterize the impacts of swapping on a range of statistical quantities of interest. We provide intuition for the types of shifts induced by swapping and compare against those introduced by TopDown. We find that even when swapping and TopDown introduce errors of similar magnitude, the direction in which statistics are biased need not be the same across the two techniques. More broadly, our implementation provides researchers with the tools to analyze and potentially correct for the impacts of disclosure avoidance systems on the quantities they study.
@@ 1 Introduction
Census statistics are used for a wide variety of planning, administrative, and redistricting purposes. In order to meet its dual burdens of providing useful statistics and protecting the privacy of individual respondents, the US Census Bureau has for decades introduced noise into published statistics. From 1990–2010, the decennial census used a form of “swapping” (Dalenius and Reiss, 1982; McKenna, 2019). The details of the method are kept confidential.
@@
At a high level, swapping works by choosing households that are at a particularly high disclosure risk—say, a household that is racially unique in its neighborhood—and “swapping” each such household with another household in a different region, “hiding” the unique household in an area where it may be less identifiable. The precise instantiation of this framework used by the Bureau is kept secret, and indeed, we do not even know whether any single algorithm was used across all states in the 1990–2010 decennial censuses.
@@
Prior to 2020, although it was generally known that swapping was deployed, consumers of Census statistics tended to treat published data as ground truth. This became impossible with the announcement of the deployment of differential privacy, the publication of demonstration data and, eventually, the TopDown algorithm itself. [^2] Differential privacy does not rely on secrecy of the algorithm to maintain privacy for respondents. In a nutshell, consumers were forced to confront the “ground truth” fiction. Each DAS, swapping and TopDown, has both privacy properties and impacts on data analysis. We know that the Bureau chose TopDown because it can provide better privacy protection than that obtained with the level of swapping used in the 2010 census. In this work, we turn to the impact of these disclosure avoidance systems on data analysis.
@@
Studying the impacts of disclosure avoidance systems is challenging due to a lack of both data and information. In the case of TopDown, the Census Bureau has provided the details of the algorithm and a series of “before and after” demonstration data releases to facilitate research (see Section 2 for details). No analogous releases exist for swapping. In this work, we facilitate research into the effects of swapping through a careful implementation of a swapping algorithm based on information released by the Census Bureau (see Section 3); specifically, we drawn from a variety of publications, presentations, and statements by Bureau employees, as well as court documents, to create an approximation to the Bureau’s practices. For details that remain unknown, we conduct robustness checks to verify that our conclusions hold across a range of reasonable choices.
@@
We use this implementation to study a variety of impacts that swapping has on the Census data. In Section 4, to investigate the basic statistical properties of swapping, we measure the errors that swapping introduces to population counts by race, and compare them to those of TopDown and a simple variant, ToyDown, created by Cohen et al. (2021), at a specific setting and apportioning of the privacy parameter $\varepsilon$. We further examine detailed properties of swapping with respect to household size, mixed-race households, and racial entropy across geographies. Surprisingly, despite the fact that swapping targets households that are racially unique in their current location and “moves” them elsewhere, we find that swapping increases racial entropy. We provide a detailed analysis of why this is the case.
@@
We next consider the impact of swapping on downstream uses of Census data for both public policy and social science (Section 5). Prior work has found that differential privacy biases statistics derived from ecological regression (ER) analyses, which are used in cases involving the Voting Rights Act (Cohen et al., 2021). We find that the same is true for swapping, but, crucially, swapping appears to bias these statistics in the opposite direction. We again provide intuition for why this might be the case. Thus, viewing swapping as a similarly noisy predecessor to differential privacy may obscure key differences in the direction of effects of that noise.
@@
We further compare swapping to TopDown in terms of impact on social science applications, where recent work has shown that the noise introduced by the TopDown algorithm can lead to substantial variance in downstream analyses (Mueller and Santos-Lozada, 2022; Hauer and Santos-Lozada, 2021; Ruggles et al., 2019; Santos-Lozada et al., 2020; Winkler et al., 2021). In particular, Mueller and Santos-Lozada (2022) show that TopDown can introduce large errors in rural areas. We find that the same is true for swapping. We conclude by proposing a more general framework for using our methods to estimate the impacts of swapping on downstream statistics of interest.
@@ 2 Related Work
Our focus in this work is on the Decennial Census. Disclosure avoidance is also used in the American Community Survey (ACS), which collects more detailed information on a sample of the population (see Zayatz et al. (2009)). The Census Bureau uses a variety of disclosure avoidance techniques to release ACS data, including swapping (Poehler, 2022, Section 13.8), (Census Bureau, 2022; Rodríguez, 2023).
@@ 3 Swapping Implementation
The swap rate is not public, but we have some information about reasonable ranges. Bailie et al. (2023) claim that the swap rate is 2–4%. John Abowd, the former Chief Scientist at the U.S. Census Bureau, states that the swap rates that the Bureau considered when studying the feasibility of using swapping in 2020 were higher than the swap rates used in 2010 (Abowd, 2021). An example of the swap rates that the Bureau considered for 2020 appears in Hawes and Rodríguez (2021) (also written by Bureau officials), in which the authors present swap rates in the range of 5–50%. Based on this information, we use 2% as a reasonable low swap rate and and 10% as a reasonable high swap rate. In some places, we also present results for higher swap rates.
@@
Documentation on matching variables is inconsistent. According to Zayatz (2002), “The key variables were number of people in the household of each race by Hispanic/Non-Hispanic by age group (<18,18+), number of units in building, rent/value of home, and tenure.” Griffin et al. (1989) states that the planned swapping methodology matches on “number of persons in household, population characteristics of race, Hispanic origin and age, and on housing characteristics of units at building, rent/value and tenure.” If true, however, this would guarantee that block-level counts by race and ethnicity are invariant to swapping, which goes against assumptions commonly held in the literature (e.g.,  Christ et al., 2022; Kenny et al., 2023b). Moreover, Steel and Zayatz (2003) appear to contradict this in their evaluation of the impacts of swapping, writing “Our procedure targeted households with unique characteristics, and this increased the number of zero cells (for these characteristics) because it tended to draw people from cells with values 1, 2, and 3 (presumably where the entire contribution to the cell is from one household and hence selected for high disclosure risk) and swap them into cells with larger values.”
@@ 4.1 Setup and notation
By running ToyDown, the simplified version of TopDown implemented by Cohen et al. (2021), on our synthetic data, we can also produce $\mathbf{X}_{\text{ToyDown}}$. To produce this database, we run ToyDown with $\epsilon=3.26$ and an equal split of privacy budget across geographic levels. This value of $\epsilon$ was chosen by calibrating the variance of ToyDown against the variance of TopDown in demonstration data (see Section 4.3 for details). Technically, our implementation of ToyDown differs from the one used in (Cohen et al., 2021) because we do not treat Hispanic as a race category to be consistent with our swapping algorithm.
@@
Figure 1: Various versions of the Census data and the relationships between them. Note that there are two releases of $\mathbf{X}_{\textnormal{TopDown}}$, which we call $\mathbf{X}_{\textnormal{TopDown2021}}$ and $\mathbf{X}_{\textnormal{TopDown2023}}$. We also produce multiple versions of $\mathbf{X}_{\textnormal{Swapped}}$: $\mathbf{X}_{\textnormal{Swapped2}}$ is produced by swapping at a 2% swap rate and $\mathbf{X}_{\textnormal{Swapped10}}$ is produced by swapping at a 10% swap rate. The red arrows show the comparisons made in the panels in Figures 2 and 3. The blue arrow shows the comparison of interest (which we cannot directly observe).
@@ 4.2 Quantifying the Errors Introduced by Swapping
We begin by studying the bias of swapping on race counts. It is important to remember that the state-wide average bias of the count of any race is zero, since swapping (unlike TopDown) does not change the statewide counts of each race. Therefore, we study the error in the individual counts at lower geographic levels. These errors will average to zero across the entire state, but we are interested in their magnitudes.
@@
Figure 2 shows $\textsf{Error}_{\gamma}^{r}(\mathbf{X}_{\text{Synthetic}},\mathbf{X}_{\text{Swapped2}})$, $\textsf{Error}_{\gamma}^{r}(\mathbf{X}_{\text{Synthetic}},$ $\mathbf{X}_{\text{Swapped10}})$, $\textsf{Error}_{\gamma}^{r}(\mathbf{X}_{\text{Synthetic}},\mathbf{X}_{\text{ToyDown}})$, and $\textsf{Error}_{\gamma}^{r}(\mathbf{X}_{\text{Released}},$
$\mathbf{X}_{\text{TopDown2021}})$ at the county level in Alabama. As noted earlier, the errors for swapping (at both 2% and 10%) are both mean zero. [^5] The swap errors do not necessarily have median zero; the blue lines in the boxplots are the medians. We observe that errors due to swapping (Panels A and B) are generally greater than errors due to ToyDown (Panel C). Also, swapping results in higher errors on large racial groups, while ToyDown distributes errors evenly across racial groups. The picture for TopDown (Panel D) may superimpose the errors due to swapping and TopDown, which is consistent with the fact that both swapping and TopDown stand between $\mathbf{X}_{\text{TopDown}}$ and $\mathbf{X}_{\text{Released}}$.
@@ C Refer to caption Refer to caption D
Figure 2: Plots of the four quantities $\textnormal{{Error}}_{\gamma}^{r}(\mathbf{X}_{\textnormal{Synthetic}},\mathbf{X}_{\textnormal{Swapped2}})$, $\textnormal{{Error}}_{\gamma}^{r}(\mathbf{X}_{\textnormal{Synthetic}},\mathbf{X}_{\textnormal{Swapped10}})$, $\textnormal{{Error}}_{\gamma}^{r}(\mathbf{X}_{\textnormal{Synthetic}},\mathbf{X}_{\textnormal{ToyDown}})$, $\textnormal{{Error}}_{\gamma}^{r}(\mathbf{X}_{\textnormal{Released}},\mathbf{X}_{\textnormal{TopDown2021}})$ for all $\gamma\in\mathcal{C}_{\textnormal{Alabama}}$, where $\mathcal{C}_{\textnormal{Alabama}}$ is the set of all counties in Alabama. To generate the “Average” plot, for each county, the errors for each race were averaged. Note that the scale of the $y$-axes are different in different plots.
@@
as a measure of the effect of swapping on the count of race $r$ in region $\gamma$, normalized by the size of the count. If the count increases by a relatively large amount from $\mathbf{X}_{1}$ to $\mathbf{X}_{2}$, $\textsf{RelativeError}_{\gamma}^{r}(\mathbf{X}_{1},\mathbf{X}_{2})$ will be close to 2. If the count decreases by a relatively large amount, $\textsf{RelativeError}_{\gamma}^{r}(\mathbf{X}_{1},\mathbf{X}_{2})$ will be close to 0. If the count is unchanged, $\textsf{RelativeError}_{\gamma}^{r}(\mathbf{X}_{1},\mathbf{X}_{2})=1$. If both counts are 0, we treat the relative error as 1. In practice, the way we treat zeros is not very important because the only county-level race counts that are zero are the number of Hawaiian or Pacific Islander people in five counties (before swapping).
Again, our ideal comparison is $\textsf{RelativeError}_{\gamma}^{r}(\mathbf{X}_{\text{CEF}},\mathbf{X}_{\text{Released}})$. However, since we do not observe $\mathbf{X}_{\text{CEF}}$, we instead measure
$\textsf{RelativeError}_{\gamma}^{r}(\mathbf{X}_{\text{Synthetic}},\mathbf{X}_{\text{Swapped}})$ (defined for swapping at a 2% swap rate and a 10% swap rate), as well as $\textsf{RelativeError}_{\gamma}^{r}(\mathbf{X}_{\text{Synthetic}},$ $\mathbf{X}_{\text{ToyDown}})$ and $\textsf{RelativeError}_{\gamma}^{r}(\mathbf{X}_{\text{Released}},\mathbf{X}_{\text{TopDown2021}})$. These are shown in Figure 3. We observe that 2% swapping has smaller relative errors than ToyDown, but 10% swapping has comparable relative errors to ToyDown. Both swapping and ToyDown exhibit very low relative error on large racial groups and much larger relative error on small racial groups (especially with a 10% swap rate, but often with a 2% swap rate as well). In the case of ToyDown, this is consistent with the fact that ToyDown distributes absolute errors relatively evenly across racial groups, so groups with smaller sizes have larger relative errors. In the case of swapping, it appears that, while swapping has larger absolute errors on larger racial groups, this is more than canceled out by the effect of the group sizes. The overall effect is that swapping, at least at a 10% swap rate, introduces larger relative errors to smaller groups.
@@
Figure 3: Plots of the four quantities $\textnormal{{RelativeError}}_{\gamma}^{r}(\mathbf{X}_{\textnormal{Synthetic}},\mathbf{X}_{\textnormal{Swapped2}})$, $\textnormal{{RelativeError}}_{\gamma}^{r}(\mathbf{X}_{\textnormal{Synthetic}},\mathbf{X}_{\textnormal{Swapped10}})$, $\textnormal{{RelativeError}}_{\gamma}^{r}(\mathbf{X}_{\textnormal{Synthetic}},\mathbf{X}_{\textnormal{ToyDown}})$, and $\textnormal{{RelativeError}}_{\gamma}^{r}(\mathbf{X}_{\textnormal{Released}},\mathbf{X}_{\textnormal{TopDown2021}})$ for all $\gamma\in\mathcal{C}_{\textnormal{Alabama}}$, where $\mathcal{C}_{\textnormal{Alabama}}$ is the set of all counties in Alabama. To generate the “Average” plot, for each county, the relative errors for each race were averaged.